# Threats, Attacks, and Defenses in Machine Unlearning: A Survey

### v1

@@ Abstract
Recently, Machine Unlearning (MU) has gained considerable attention for its potential to improve AI safety by removing the influence of specific data from trained Machine Learning (ML) models. This process, known as knowledge removal, addresses concerns about data such as sensitivity, copyright restrictions, obsolescence, or low quality. This capability is also crucial for ensuring compliance with privacy regulations such as the Right To Be Forgotten (RTBF). Therefore, strategic knowledge removal mitigates the risk of harmful outcomes, safeguarding against biases, misinformation, and unauthorized data exploitation, thereby enhancing the ethical use and reliability of AI systems. Efforts have been made to design efficient unlearning approaches, with MU services being examined for integration with existing machine learning as a service (MLaaS), allowing users to submit requests to erase data. However, recent research highlights vulnerabilities in machine unlearning systems, such as information leakage and malicious unlearning requests, that can lead to significant security and privacy concerns. Moreover, extensive research indicates that unlearning methods and prevalent attacks fulfill diverse roles within MU systems. For instance, unlearning can act as a mechanism to recover models from backdoor attacks, while backdoor attacks themselves can serve as an evaluation metric for unlearning effectiveness. This underscores the intricate relationship and complex interplay between these elements in maintaining system functionality and safety. Therefore, this survey seeks to bridge the gap between the extensive number of studies on threats, attacks, and defenses in machine unlearning and the absence of a comprehensive review that categorizes their taxonomy, methods, and solutions, thus offering valuable insights for future research directions and practical implementations.
@@ I Introduction
  • Training data protected by copyright requires the removal of knowledge gained from them from the trained ML model. This step is crucial to comply with commercial regulations and avoid conflicts before the model’s deployment or release.
  • Additionally, global data privacy laws like GDPR [1], APPI [2], and CCPA [3] emphasize data rights, including the Right To Be Forgotten (RTBF), allowing individuals to request the deletion of their personal data. This underscores the importance of effective knowledge removal to comply with these regulations.
@@
As knowledge removal in AI systems necessitates the elimination of certain data’s presence and influence from a trained ML model, machine unlearning (MU) [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] emerges as a crucial enabler of this process. In a MU system, the training dataset $D$ comprises two subsets, namely $D_{u}$ and $D_{r}$, representing the data points to be unlearned and the remaining data points, respectively, wherein $D_{r}=D\backslash D_{u}$. Denote $\mathcal{M}_{t}=\mathcal{M}(D)$ as the model trained on the entire dataset $D$. The goal of MU is to produce an unlearned model $\mathcal{M}_{u}$ that closely resembles the model $\mathcal{M}_{u}^{*}$ trained solely on $D_{r}$, i.e., $\mathcal{M}_{u}^{*}=\mathcal{M}(D_{r})$.
As depicted in Figure 1, a naive yet initial approach to implementing MU is through retraining. This method involves discarding the current model and retraining it from scratch with the remaining data after the removal of the data to be unlearned. Nevertheless, this approach can incur substantial computational expenses, making it impractical for scenarios involving large-scale models or extensive datasets. In response to these challenges, a range of efficient MU methodologies have recently been developed, focusing on either exact or approximate unlearning [17]. Exact unlearning techniques typically involve retraining but limit the scope of data involved to enhance efficiency over naive retraining approaches. On the other hand, approximate unlearning adjusts the parameters of the existing trained model to create an unlearned model that approximates one that would be obtained through retraining from scratch. This approach often results in a tradeoff between unlearning performance and efficiency. We recommend interested readers consult survey papers on machine unlearning, such as [17, 9, 10, 19, 20], for more comprehensive details about MU methodologies and taxonomy. For better understanding, we first provide some definitions and previews in the subsequent part, while our contributions are also clarified.
@@ I-A Machine Unlearning Systems
Note that these roles can be allocated to different entities depending on the application scenario. In addition, the roles of participants within a MU system are classified as follows:
@@
Note that akin to the server’s multi-functional role, participants within unlearning systems can embody various roles. For instance, it is common for data contributors to also have the authority to submit unlearning requests, icompliance with RTBF regulations. Additionally, the workflow of machine unlearning systems can be structured into three phases, which we categorize as follows:
@@
  • Unlearning phase: the server eliminates certain data points from the model’s knowledge base, in response to unlearning requests from requesting users.
@@ I-B Threats, Attacks, and Defenses in MU Systems
Besides threats and attacks on machine unlearning systems, research shows that unlearning and attacks serve dual roles. This reveals a complex interaction between these elements, essential for system safety and functionality. In this section, we offer a comprehensive overview of threats, attacks, and defenses in machine unlearning systems. We explore these topics in depth in Sections III, IV, and V, where we present a detailed taxonomy as shown in Figure 3.
@@
Threats. As illustrated in Figure 2, vulnerabilities are present across all phases of the unlearning system. In the training phase, for example, data contributors can maliciously craft the training data in a way that prepares for future attacks. In the unlearning phase, requesting users have the opportunity to submit malicious unlearning requests that could be designed to further their goals via the unlearning mechanism. Moreover, during the post-unlearning phase, participants with access to both the trained and the unlearned models might exploit this dual access to uncover additional information leakages, thus exposing further vulnerabilities created by the unlearning process.
Attacks. The threats outlined previously can be exploited to execute attacks. For instance, a data contributor might poison the training dataset, and bypass the detection by incorporating mitigation data during the training phase, which could then be removed via unlearning in the unlearning phase [21]. A requesting user could craft an unlearning request to unlearn more information than expected, thereby manipulating the system to their advantage [22]. Furthermore, an accessible user could perform enhanced membership inference attacks [23, 24, 25] by leveraging their access to both the trained and unlearned models, exploiting the differential insights provided by comparing the two models [26].
Beyond the attacks on machine unlearning systems, extensive research has demonstrated that certain prevalent attack strategies, like backdoor attacks [27] and membership inference attacks [23], can be adapted for unlearning evaluations, such as privacy leakage audit [26], model robustness assessment [22], and proof of unlearning [28].
Defenses. Potential defense methods against attacks on MU systems are explored in conjunction with attack strategies. For example, performing membership checks on data targeted for unlearning can ensure that the data specified in unlearning requests genuinely exists in the training dataset [22]. Implementing differential privacy [29, 30] can help prevent membership inference attacks, with a sacrifice of model performance [26]. Moreover, anomaly detection through model monitoring in the unlearning phase can effectively identify potential issues [31].
Additionally, unlearning can serve as a method of defense. It is capable of removing backdoors from trained machine learning models [32, 33, 34], and can be employed to achieve value alignment, thereby providing a defense mechanism against potential AI safety issues [35, 36, 37].
@@ I-C Our Contributions
With the growing focus on security and privacy in unlearning systems, the vulnerabilities within these systems and the threats they face have become increasingly apparent. Furthermore, the multifaceted roles of unlearning methods and common attacks are becoming increasingly apparent, emphasizing their complex interplay in protecting the functionality and safety of MU systems. Given this backdrop, a thorough examination of existing research on threats, attacks, and defenses in MU systems is indispensable. Such a comprehensive review would not only categorize their taxonomy, methods, and solutions but also provide valuable insights for future research directions and practical implementations. Consequently, this survey is developed to address this existing gap, aiming to offer a holistic overview and critical analysis of the field.
Comparison with other surveys. Recently, several studies have focused on surveying and analyzing machine unlearning methodologies, resulting in a clear categorization of the existing survey research. These categories include general surveys on machine unlearning, highlighted by works such as [8, 17, 9, 15, 10], which cover a wide range of topics from basic concepts and methods to the challenges faced in the field. Additionally, specific attention has been given to federated unlearning (FU) through studies like [19, 38, 39, 20]. These surveys focus on the unique issues and uses of federated unlearning compared to traditional unlearning approaches. Furthermore, potential risks and threats within federated unlearning are examined in [40].
To the best of our knowledge, there currently exists no comprehensive survey that specifically tackles the issues of security and privacy within the context of machine unlearning, particularly focusing on threats, attacks, and defensive strategies. This gap in the literature strongly motivates our work in compiling this survey. Our focus on categorizing and scrutinizing these elements aims to illuminate strategies for protecting machine unlearning systems, thereby enhancing their robustness and safety.
@@
    We present a unified workflow for machine unlearning, and drawing from this framework as well as different threat models, we propose a novel taxonomy of threats, attacks, and defenses within MU systems.
@@
    We delve deeply into critical discussions concerning the potential threats, attacks, and defenses within machine unlearning, highlighting key areas for future research exploration.
@@ III-A Information Leakage
Note that the privacy leakages mentioned above occur in the post-unlearning phase when accessible users can interact with the model service for inferences. However, [42] highlights a new privacy risk from the Rignt-To-Be-Forgetten perspective in a machine-learning-as-a-service (MLaaS) setting where the unlearning phase for one request and the post-unlearning phase for another inference request may overlap, i.e., inference requests and unlearning requests may arrive at the server within a very close timeframe. In such scenarios, prioritizing the “unlearning-request-first” could lead to service obsolescence, whereas handling the “inference-request-first” enables an attacker to gain knowledge about a specific input, which can be exploited by attacks like membership inference attacks. To counter this threat, strategies have been developed to evaluate the urgency of processing unlearning requests immediately. This evaluation is based on whether the unlearning disrupts the consistency between the trained and unlearned models.
@@ III-C Other Vulnerabilities
Threats Ref.
R1
R2
R3
G1
G2
G3
G4
K1
K2
K3
P1
P2
P3
Techniques Defenses
[26] Membership inference attacks
Limiting output knowledge (S3); Label smoothing (S3); Differential privacy (S3)
[41]
Membership inference attacks
Label smoothing (S3); Differential privacy (S3)
[43]
Deletion inference attacks; Deletion reconstruction attacks
Differential privacy (S3)
Model Discrepancy [42]
Obsolescent inference services
Strategic postponed unlearning (S2,S3)
[44]
Submitting adaptive requests; Storing secret models
Regulated unlearning algorithms (S1); Redefined certified removal (S1)
Information Leakage Knowledge Dependency [45]
Adversarial embeddings
NA
[22] Over-unlearning
Membership checking (S1); Model monitoring (S2)
Direct Unlearning Attacks [47]
Perturbing unlearning requests
NA
@@
[31]
Camouflaged attacks
Model monitoring (S2)
@@
[53]
Poisoning training data
NA
@@
Other Vulnerabilities [58]
Over-forgetting
Improved unlearning algorithms (S1)
@@ IV-A Model Recovery
In this context, machine unlearning presents itself as an intuitive approach for model recovery, enabling the targeted removal of identified poisoned data. For example, [34] demonstrates backdoor removal by inverting the trigger to retrain infected ML models, a technique also utilized in [69]. A more complex strategy for eliminating backdoors is detailed in [32], where the process begins with identifying the trigger pattern based on which then conducting unlearning via gradient ascent. In [37], sparsity-aware unlearning is executed by initially pruning the model, followed by the unlearning process. Diverging from these approaches, [70] and [71] explore the effects of unlearning on adversarial and backdoor attacks, each proposing different methods to balance trade-offs. The former focuses on unlearning universal adversarial perturbations, while the latter targets the unlearning of shared adversarial examples. The concept of model recovery through unlearning has also been explored in various other settings, including multimodal environments as seen in [72], and federated learning scenarios, as discussed in [33, 73].
@@ V-B Assessment of Model Robustness
Similarly, malicious unlearning can assess the robustness of the unlearned model. The foundational principle of this assessment approach is akin to that used in adversarial attacks, as detailed in works such as [85, 86, 87, 88, 89]. For example, [22] describes a method of malicious unlearning via over-unlearning, where perturbed data are submitted in the unlearning request to alter the classifier’s decision boundary. This strategy of pushing the decision boundary allows for evaluating the impact of unlearned data points near the boundary on model performance, as well as the effect of data points that challenge the model’s prediction capabilities. Another example occurs in the context of LLMs[45], where attackers insert adversarial embeddings into the prompts’ embeddings, aiming to extract supposedly unlearned knowledge from the model. Such attacks provide a means to gauge the model’s resilience against malicious input embeddings. On a different note, [90] and [91] treat data removal as an adversarial perturbation to the overall training dataset. In this case, the concept of randomized smoothing [92] can be adopted to measure the model robustness against unlearned data [90] or unlearned users in a federated setting [91], hence establishing a certified budget for data removals, within which the model retains its robustness.

### v2

@@ Abstract
Machine Unlearning (MU) has gained considerable attention recently for its potential to achieve Safe AI by removing the influence of specific data from trained Machine Learning (ML) models. This process, known as knowledge removal, addresses AI governance concerns of training data such as quality, sensitivity, copyright restrictions, and obsolescence. This capability is also crucial for ensuring compliance with privacy regulations such as the Right To Be Forgotten (RTBF). Furthermore, effective knowledge removal mitigates the risk of harmful outcomes, safeguarding against biases, misinformation, and unauthorized data exploitation, thereby enhancing the safe and responsible use of AI systems. Efforts have been made to design efficient unlearning approaches, with MU services being examined for integration with existing machine learning as a service (MLaaS), allowing users to submit requests to remove specific data from the training corpus. However, recent research highlights vulnerabilities in machine unlearning systems, such as information leakage and malicious unlearning requests, that can lead to significant security and privacy concerns. Moreover, extensive research indicates that unlearning methods and prevalent attacks fulfill diverse roles within MU systems. For instance, unlearning can act as a mechanism to recover models from backdoor attacks, while backdoor attacks themselves can serve as an evaluation metric for unlearning effectiveness. This underscores the intricate relationship and complex interplay among these mechanisms in maintaining system functionality and safety. This survey aims to fill the gap between the extensive number of studies on threats, attacks, and defenses in machine unlearning and the absence of a comprehensive review that categorizes their taxonomy, methods, and solutions, thus offering valuable insights for future research directions and practical implementations.
@@ I Introduction
  • Training data protected by copyright requires the removal of knowledge gained from them from the trained machine learning (ML) model. This step is crucial to comply with commercial regulations and avoid conflicts before the model’s deployment or release.
  • Additionally, data privacy regulations such as GDPR [1], APPI [2], and CCPA [3] emphasize data rights, including the Right To Be Forgotten (RTBF), allowing individuals to request the deletion of their personal data. This underscores the importance of effective knowledge removal to comply with these regulations.
@@
In AI systems, removing knowledge requires the deletion of specific data and its effects from a trained ML model. As such, machine unlearning (MU) [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] emerged as a crucial enabler of this process. In a MU system, the training dataset $D$ comprises two subsets, namely $D_{u}$ and $D_{r}$, representing the data points to be unlearned and the remaining data points, respectively, wherein $D_{r}=D\backslash D_{u}$. Denote $\mathcal{M}_{t}=\mathcal{M}(D)$ as the model trained on the entire dataset $D$. The goal of MU is to produce an unlearned model $\mathcal{M}_{u}$ that closely resembles the model $\mathcal{M}_{u}^{*}$ trained solely on $D_{r}$, i.e., $\mathcal{M}_{u}^{*}=\mathcal{M}(D_{r})$.
As depicted in Figure 1, an initial, yet naive, approach to implementing MU is through retraining. This method involves discarding the current model and retraining it from scratch with the remaining data after the removal of the data to be unlearned. Nevertheless, this approach can incur substantial computational expenses, making it impractical for scenarios involving large-scale models or extensive datasets. In response to these challenges, a range of efficient MU methodologies have recently been developed, focusing on either exact or approximate unlearning [18]. Exact unlearning techniques typically involve retraining but limit the scope of data involved in order to enhance efficiency compared to naive retraining approaches. On the other hand, approximate unlearning adjusts the parameters of the existing trained model to create an unlearned model that approximates one that would be obtained through retraining from scratch. This approach often requires tradeoffs between unlearning performance and efficiency. Interested readers are referred to the relevant survey papers on machine unlearning, such as [18, 9, 10, 20, 21, 22], for more comprehensive details about MU methodologies and taxonomy. For better understanding, we first provide some definitions and previews in the subsequent part, while our contributions are also clarified.
@@ I-A Machine Unlearning Systems
Note that these roles can be allocated to different entities depending on the application scenario [23]. In addition, the roles of participants within a MU system are classified as follows:
@@
Note that akin to the server’s multi-functional role, participants within unlearning systems can embody various roles. For instance, it is common for data contributors to also have the authority to submit unlearning requests, in compliance with RTBF regulations. Additionally, the workflow of machine unlearning systems can be structured into three phases, which we categorize as follows:
@@
  • Unlearning phase: the server conducts unlearning algorithms to remove the effects of certain data points from the model’s knowledge base, in response to unlearning requests from requesting users.
@@
In the context of machine-learning-as-a-service (MLaaS), the server may receive requests for unlearning and inference simultaneously. This indicates a possible overlap between the unlearning phase and the post-unlearning phase, where the procedures for managing different types of requests become more complex [24].
@@ I-B Threats, Attacks, and Defenses in MU Systems
Vulnerabilities are present in machine unlearning systems, where adversaries can launch various attacks at different phases, posing significant threats to the safe deployment of unlearning services within MLaaS. Besides threats and attacks on machine unlearning systems, many studies show that unlearning and attacks serve dual roles. For instance, unlearning can act as a mechanism to recover models from backdoor attacks, while backdoor attacks themselves can serve as an evaluation metric for unlearning effectiveness. This reveals a complex interaction between these elements, essential for system safety and functionality. In this section, we offer an overview of threats, attacks, and defenses in machine unlearning systems. We explore these topics in depth in Sections III, IV, and V, where we present a detailed taxonomy as shown in Figure 3.
@@
Threats. As illustrated in Figure 2, vulnerabilities are present across all phases of the unlearning system. In the training phase, for example, data contributors can maliciously craft the training data in a way that prepares for future attacks. In the unlearning phase, requesting users have the opportunity to submit malicious unlearning requests that could be designed to achieve their goals after the unlearning process. Moreover, during the post-unlearning phase, participants with access to both the trained and the unlearned models might exploit this dual access to uncover additional information leakages, thus exposing further vulnerabilities created by the unlearning process. A more detailed discussion on the various threats that may arise is available in Section III.
Attacks. The threats outlined previously can be exploited to execute attacks. For instance, a data contributor might poison the training dataset, and bypass the detection by incorporating mitigation data during the training phase, which could then be removed via unlearning in the unlearning phase [25]. A requesting user could craft an unlearning request to unlearn more information than expected, thereby manipulating the system to their advantage [23]. Furthermore, an accessible user could perform enhanced membership inference attacks [26, 27, 28] by leveraging their access to both the trained and unlearned models, exploiting the differential insights provided by comparing the two models [29]. These attacks, along with threats to unlearning systems, are discussed in detail in Section III.
Beyond the attacks on machine unlearning systems, extensive research has demonstrated that certain prevalent attack strategies, like backdoor attacks [30] and membership inference attacks [26], can be adapted for unlearning evaluations, such as privacy leakage audit [29], model robustness assessment [23], and proof of unlearning [31]. The analysis of attacks used for evaluating unlearning is provided in Section IV.
Defenses. Potential defense methods against attacks on MU systems are explored in conjunction with attack strategies. For example, performing membership checks on data targeted for unlearning can ensure that the data specified in unlearning requests genuinely exists in the training dataset [23]. Implementing differential privacy [32, 33] can help prevent membership inference attacks, with a sacrifice of model performance [29]. Moreover, anomaly detection through model monitoring in the unlearning phase can effectively identify potential issues [34]. A corresponding discussion, along with an analysis of threats and attacks, is given in Section III.
Additionally, unlearning can serve as a method of defense. It is capable of removing backdoors from trained machine learning models [35, 36, 37], and can be employed to achieve value alignment, thereby providing a defense mechanism against potential AI safety issues [38, 39, 40]. The dual role of unlearning as a defense mechanism is analyzed in Section IV.
@@ I-C Our Contributions
The increasing emphasis on security and privacy in unlearning systems has led to a clearer recognition of their vulnerabilities and the threats they face, as revealed by related studies. Moreover, the complex roles of unlearning methods and the prevalence of attacks have become more evident, underscoring their intricate interaction in maintaining the functionality and safety of MU systems. Given this backdrop, a thorough examination of existing research on threats, attacks, and defenses in MU systems is indispensable. Such a comprehensive review would not only categorize their taxonomy, methods, and solutions but also provide valuable insights for future research directions and practical implementations. Consequently, this survey is developed to address this existing gap, aiming to offer a holistic overview and critical analysis of the field.
Comparison with other surveys. Recently, several studies have focused on surveying and analyzing machine unlearning methodologies, resulting in a clear categorization of the existing survey research. These categories include general surveys on machine unlearning, highlighted by works such as [8, 18, 9, 16, 41, 10], which cover a wide range of topics from basic concepts and methods to the challenges faced in the field. Additionally, specific attention has been given to federated unlearning (FU) through studies like [20, 42, 43, 21]. These surveys focus on the unique issues and uses of federated unlearning compared to traditional unlearning approaches. Furthermore, potential risks and threats within federated unlearning are examined in [44]. Benchmarks and surveys on MU approaches over large language models (LLMs) are given in [45, 46, 47, 48, 39].
  • General surveys on MU: [8, 18, 9, 16, 10, 41]
  • Surveys on FU: [20, 42, 43, 21, 44]
  • Benchmarks and surveys on MU over LLMs: [45, 46, 47, 39, 48]
To the best of our knowledge, no comprehensive survey currently tackles security and privacy issues within the context of machine unlearning, particularly focusing on threats, attacks, and defensive strategies. This gap in the literature strongly motivates our work in compiling this survey. Our focus on categorizing and scrutinizing these elements aims to illuminate strategies for protecting machine unlearning systems, thereby enhancing their robustness and safety.
@@
    We present a unified workflow for machine unlearning, and drawing from this framework as well as threat models, we propose a novel taxonomy of threats, attacks, and defenses within MU systems.
@@
    We delve deeply into critical discussions concerning potential vulnerabilities and threats in unlearning systems and the challenges in defenses against these threats and attacks, highlighting key areas for future research exploration.
@@ III-A Information Leakage
Note that the privacy leakages mentioned above occur in the post-unlearning phase when accessible users can interact with the model service for inferences. However, [24] highlights a new privacy risk from the Rignt-To-Be-Forgetten perspective in an MLaaS setting where the unlearning phase for one request and the post-unlearning phase for another inference request may overlap, i.e., inference requests and unlearning requests may arrive at the server within a very close timeframe. In such scenarios, prioritizing the “unlearning-request-first” could lead to service obsolescence, whereas handling the “inference-request-first” enables an attacker to gain knowledge about a specific input, which can be exploited by attacks like membership inference attacks. To counter this threat, strategies have been developed to evaluate the urgency of processing unlearning requests immediately. This evaluation is based on whether the unlearning disrupts the consistency between the trained and unlearned models.
@@ III-C Other Vulnerabilities
Threats Ref. R1 R2 R3 G1 G2 G3 G4 K1 K2 K3 P1 P2 P3 Techniques Defenses
[29] Membership inference attacks Limiting output knowledge (S3); Label smoothing (S3); Differential privacy (S3)
[49] Membership inference attacks Label smoothing (S3); Differential privacy (S3)
[50] Deletion inference attacks; Deletion reconstruction attacks Differential privacy (S3)
Model Discrepancy [24] Obsolescent inference services Strategic postponed unlearning (S2,S3)
[51] Submitting adaptive requests; Storing secret models Regulated unlearning algorithms (S1); Redefined certified removal (S1)
Information Leakage Knowledge Dependency [52] Adversarial embeddings NA
[23] Over-unlearning Membership checking (S1); Model monitoring (S2)
Direct Unlearning Attacks [54] Perturbing unlearning requests NA
@@
[34] Camouflaged attacks Model monitoring (S2)
@@
[60] Poisoning training data NA
@@
Other Vulnerabilities [65] Over-forgetting Improved unlearning algorithms (S1)
@@ IV-A Model Recovery
In this context, machine unlearning presents itself as an intuitive approach for model recovery, enabling the targeted removal of identified poisoned data. For example, [37] demonstrates backdoor removal by inverting the trigger to retrain infected ML models, a technique also utilized in [76]. A more complex strategy for eliminating backdoors is detailed in [35], where the process begins with identifying the trigger pattern based on which then conducting unlearning via gradient ascent. [77] focuses on model recovery through unlearning, based solely on limited access to poisoned data. In [40], sparsity-aware unlearning is executed by initially pruning the model, followed by the unlearning process. Diverging from these approaches, [78] and [79] explore the effects of unlearning on adversarial and backdoor attacks, each proposing different methods to balance trade-offs. The former focuses on unlearning universal adversarial perturbations, while the latter targets the unlearning of shared adversarial examples. The concept of model recovery through unlearning has also been explored in various other settings, including multimodal environments as seen in [80], and federated learning scenarios, as discussed in [36, 81].
@@ V-B Assessment of Model Robustness
Similarly, malicious unlearning can assess the robustness of the unlearned model. The foundational principle of this assessment approach is akin to that used in adversarial attacks, as detailed in works such as [93, 94, 95, 96]. For example, [23] describes a method of malicious unlearning via over-unlearning, where perturbed data are submitted in the unlearning request to alter the classifier’s decision boundary. This strategy of pushing the decision boundary allows for evaluating the impact of unlearned data points near the boundary on model performance, as well as the effect of data points that challenge the model’s prediction capabilities. Another example occurs in the context of LLMs[52], where attackers insert adversarial embeddings into the prompts’ embeddings, aiming to extract supposedly unlearned knowledge from the model. Such attacks provide a means to gauge the model’s resilience against malicious input embeddings. On a different note, [97] and [98] treat data removal as an adversarial perturbation to the overall training dataset. In this case, the concept of randomized smoothing [99] can be adopted to measure the model robustness against unlearned data [97] or unlearned users in a federated setting [98], hence establishing a certified budget for data removals, within which the model retains its robustness.